OpenStudy (anonymous):

A baseball is hit straight up and is caught by the catcher 2.0 s later. The maximum height of the ball during this interval is:

OpenStudy (raffle_snaffle):

There isn't anymore information given?

OpenStudy (anonymous):

no

OpenStudy (ineedserioushelp):

The answer is: 12.6m . I had this question before and to find the answer use this: d = (1/2)at^2 with t = 1 second and a = 9.8 m/sec^2 . It took 1 second to go in the air at its top speed and 1 second to fall down.

OpenStudy (anonymous):

the answer is 4.9

OpenStudy (anonymous):

i tried to use max height equation 1/2 V^2(initial in respect to y)/ g

ganeshie8 (ganeshie8):

4.9 is correct !

OpenStudy (anonymous):

how did you get it????

ganeshie8 (ganeshie8):

easy, since it took 2s for going up and falling down, it took 1s for going up

ganeshie8 (ganeshie8):

agree so far ?

OpenStudy (anonymous):

yes!

ganeshie8 (ganeshie8):

Next, when it reaches top most, its velocity becomes 0 so, v = u + at 0 = u -g(1) u = g = 9.8 m/s

ganeshie8 (ganeshie8):

so initial velocity = u = 9.8 m/s

OpenStudy (anonymous):

wait how is acceleration negative?

ganeshie8 (ganeshie8):

now, use the position equation :- S = ut + 1/2 at^2 S = 9.8(1) - 1/2 g (1)^2 S = 9.8 - 1/2 *9.8 S = 9.8 - 4.9

ganeshie8 (ganeshie8):

very good q

OpenStudy (anonymous):

how i'm setting things up is up is positive and down is negative

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):